the number of marbles in jar A is 4 times than in jar B.If 96 marbles are transferred from jar A to jar B,the number of marbles in jar B will become 4 times than in jar A.Find the total number of marbles in both jar?

Answer by stanbon(75887)

the number of marbles in jar A is 4 times than in jar B.If 96 marbles are transferred from jar A to jar B,the number of marbles in jar B will become 4 times than in jar A.Find the total number of marbles in both jar?
----
Equations:
A = 4B
B+96 = 4(A-96)
---
Substitute for "A" and solve for "B":
B+96 = 4(4B-96)
B + 96 = 16B-4*96
---
15B = 5*96
B = (1/3)96
B = 32 (# of marbles in jar B)
---
Solve for "A":
A = 4B
A = 4*32
A = 128 (# of marbles in A)
================================
Cheers,
Stan H.
=============